How would I write this using positive exponents?
a^-2bc^-5/a^5bc^-4
I know that if you have a negative exponent, in order to make it positive you write it as a fraction. Example a^-n =1/a^n
I have come up with the following answer, but don't think it is correct. Please help!Thank you!
1b11/a^2a^5bc^5c^4
b/a^10bc^20

Answer by mmm4444bot(95)

Hello There:
You are correct when saying that a^(-n) = 1/a^n, but I like a different approach:
a^n/a^m = a^(n - m).
In other words, if the same variable appears in both the numerator and denominator, you can simplfy it by subracting the exponent in the denominator from the exponent in the numerator.
Your expression contains the variables a, b, and c in both the numerator and denominator, so this rule applies nicely.
[a^(-2)*b*c^(-5)]/[a^5*b*c^(-4)]
a^(-2 - 5)*b^(1-1)*c^(-5 + 4)
a^(-7)*b^0*c^(-1)
a^(-7)*c^(-1)
Now we can use the rule that you typed to get the simplified expression with no negative exponents.
1/(a^7*c)
~ Mark